# Kill-Chain Canaries: Stage-Level Tracking of Prompt Injection Across Attack Surfaces and Model Safety Tiers Haochuan Kevin Wang

### 30 Mar 2026 v1

@@ 7 Limitations
Per-cell $n$ ranges from 8 to 36; the strongest findings (Claude 0/20 per scenario, DeepSeek memory 0/24, GPT-4o-mini memory 12/12) are consistent across batches but should be interpreted with appropriate uncertainty. Payloads are explicit and synthetic; real-world attacks typically use obfuscation and social engineering. Five frontier models from three providers; reasoning models (o3, o4-mini) and open-weight models (Llama 3.3) are not represented. All defenses are lightweight wrappers that approximate the stated mechanism. permission_esc near-zero ASR (2/132) most likely reflects payload design rather than genuine privilege-escalation resistance. Classifier and provenance validation use the 192-run scenario_compare dataset; revalidation on the full 764-run set is ongoing. We do not isolate whether Claude’s summarization-stage filtering arises from training data, system prompt format, or tool API design.
@@ 8 Conclusion
We introduced kill-chain canary tracking for localizing prompt injection defenses in LLM agent pipelines. Across 764 runs and five frontier models, the key structural finding is threefold: the safety gap between models concentrates at the summarization stage, not at exposure; static defenses fail not because they are poorly designed but because they cover different surfaces than the actual injection vectors; and a Claude relay node provides measurable decontamination for downstream agents regardless of those agents’ individual safety levels. Stage-level evaluation reveals that prompt injection defenses are fundamentally pipeline-local, not model-global. The minimum evaluation standard for agentic systems should therefore include: stage-level kill-chain decomposition, multi-surface injection coverage, and relay decontamination rate as a first-class metric.

### 03 Apr 2026 v2

@@ 7 Limitations
Sample size. Per-cell $n$ ranges from 3 (Phase 3) to 36 (text-surface scenarios). The strongest text-surface findings (Claude 0/20 per scenario, DeepSeek memory 0/24, GPT-4o-mini memory 12/12) are consistent across independent batches; the Phase 3 cells at $n=3$ have wide Wilson CIs (21–94% for the 67% relay result) and must be treated as pilot-scale hypotheses.
Single task. All scenarios use one task type (document summarization with an email delivery action). Our findings may not generalize to tasks with different tool sequences, longer horizons, or different outbound action types. This is the primary threat to external validity; replication across diverse task distributions is required before deployment recommendations can be made with confidence.
Canary detection confound. A model that refuses an injection but quotes the canary token in its refusal explanation would be incorrectly labeled EXE=True. We inspected all 22 compromised text-surface runs manually and found no such false positives, but this confound is possible in principle and is not systematically controlled.
Explicit payloads. Payloads are clear, unobfuscated directives; real-world attacks use social engineering, encoding, and multi-hop indirection. Surface-mismatch results are likely conservative: defenses designed for the wrong surface will also fail against obfuscated off-surface attacks.
Other limitations. Five frontier models from three providers; reasoning models (o3, o4-mini) and open-weight models (Llama 3.3, Mistral) are not evaluated. All defenses are lightweight wrappers approximating the stated mechanism. We do not isolate whether Claude’s summarization-stage filtering arises from training, system prompt, or tool API design. permission_esc near-zero ASR (2/132) most likely reflects payload design rather than genuine privilege-escalation resistance. GPT-5-mini-2025-02-15 produced 0% tool-call rate in Phase 3; we cannot determine whether this is a capability regression or API schema incompatibility from black-box evaluation.
@@ 8 Conclusion
Stage-level canary tracking across 950 runs reveals a simple but consequential structure: prompt injection is not a model-capability problem, it is a pipeline-architecture problem. Every model is fully exposed; the safety gap lives entirely in what happens to an injection after it enters context. We draw four findings and, more importantly, four prescriptive claims about how agentic systems should be built and evaluated.
Finding 1 $\to$ Claim: Write-node placement is a deployable safety primitive today.
Claude eliminates injections at write_memory summarization on every surface we tested (0/164 text-surface, 0/9 Phase 3 PDF runs as Agent A; 95% CI 0–2%). This is not a property of Claude’s overall capability level—it is a property of its write-stage behavior. The practical implication is immediate: in any multi-agent pipeline where components can be chosen, routing all inter-agent memory writes through a safety-verified write node provides pipeline-wide decontamination independent of downstream agent choices. Conversely, Block B results show Claude as Agent B propagates at 67%—so write-node placement is the only position where this protection materializes. Future agent frameworks should expose write-node identity as a first-class architectural choice, not a deployment afterthought.
Finding 2 $\to$ Claim: Defense design must be surface-aware, not assumption-aware.
All four evaluated defenses—write_filter, pi_detector, spotlighting, and their combination—fail on at least one injection surface, not because they are incorrectly implemented but because they were designed for a different surface than the one under attack. This failure mode requires no adversarial adaptation. The correct design principle is surface decomposition before defense selection: enumerate every channel through which untrusted content can enter the pipeline, then select or compose defenses that cover each channel’s specific injection point. A monolithic defense evaluated only on the surface it was designed for is not a security evaluation.
Finding 3 $\to$ Claim: Memory provenance is a missing infrastructure primitive.
The write-vs-read safety asymmetry (Claude blocks at write, propagates at read) suggests that agent memory currently lacks the infrastructure to carry trust provenance: information about which agent wrote a memory record, under what safety context, and from what source surface. If Agent B could interrogate “was this memory written by a safety-certified node from a document I trust?” it could apply calibrated skepticism to inherited content rather than treating all memory reads as equivalently trusted delegation. Implementing content-addressed, provenance-tagged memory stores is a concrete near-term architectural direction for multi-agent frameworks.
Finding 4 $\to$ Claim: Evaluation coverage is the primary security gap, not model capability.
DeepSeek’s 0%/100% ASR split across memory and tool-stream surfaces, and the audio injection null result (spoken instructions treated as narration, not directives), show that apparent security posture is determined almost entirely by which surfaces an evaluation covers. A model that achieves 0% ASR on a single surface can achieve 100% on the next. The community standard for agentic security evaluation must require: (i) kill-chain stage decomposition rather than outcome-only ASR; (ii) multi-surface injection coverage including document extraction, tool response streams, audio transcripts, and pre-seeded memory; (iii) cross-model relay testing with heterogeneous agent pairs; and (iv) relay decontamination rate measured separately at write and read positions. Until these become baseline requirements, security evaluations of agentic systems will systematically underestimate real attack surfaces.